OpenStudy (deadshot):

If f(x) = x plus 2, all over 2, then what is f^-1(x) ?

OpenStudy (kropot72):

\[y=\frac{x+2}{2}\] The next step is to interchange x and y. Do you want me to take you thru it?

OpenStudy (anonymous):

2(x-1)

OpenStudy (anonymous):

=2(x-1)

OpenStudy (deadshot):

yes, i don't understand this kind of problem

OpenStudy (anonymous):

okay let's do it step by step

OpenStudy (kropot72):

When we interchange x and y we get \[x=\frac{y+2}{2}\] Now we need to solve for y by multiplying both sides by 2 giving\[2x=y+2\] Now you can subtract 2 from both sides to find y which is \[f ^{-1}(x)\]

OpenStudy (deadshot):

so then, it would be 2x - 2 = y, right?

OpenStudy (kropot72):

@DeadShot Good work. Youare correct. You can factorise 2x - 2 = 2(x - 1)

OpenStudy (anonymous):

f(x) = (x+2)/2 =>y=(x+2)/2 2y= x+2 x=2y-2 x=2(y-1) find f^-1(x) f^-1(x) = 2(x-1)

OpenStudy (deadshot):

Thanks!

OpenStudy (kropot72):

You're welcome :)

OpenStudy (anonymous):

you are welcome Deadshot

OpenStudy (anonymous):

so is it 2x-2 or 2x-1?